OpenStudy (anonymous):

Determine the slope and the y-intercept of the graph of the equation. x + y = 10

OpenStudy (anonymous):

y=mx + c y = 10 - x y = -x + 10 gradient: -1 y-intercept: 10

OpenStudy (radar):

I will work that out for you. The slope intercept form is y=mx+b. First isolate y, y=-x+10 by inspection we see slope m= -1, intercept = 10